1. If you buy a semi-annually compounded 5-year corporate coupon bond with a face value of $1000, coupon rate of 4%, and yield to maturity of 6%, then you know that

a)the fair price of the bond is less than $1000.

b)the coupon amount is $30.

c)both a) and b) are correct.

d)neither a) nor b) is correct.

2. Assuming 365 days in a year, if the annual interest rate is 10%, what is the present value of a $100 that you are going to receive 2-years from now with daily compounding?

a)$81.88

b)$82.64

c)$90.48

d)$122.14
Answer #1

1. Option a is Correct, Fair Price of the bond can be obtained by using the Present Value approach i.e, discounting the expected cashflows at an appropriate interest rate.

In the given question, interest rate is 4%, the PV factor as per the information provided is 0.6756, if you multiply the face value of the bond with this PV Factor, you get the fair price of the bond, i.e, 1000*0.6756=675.6

PV Factor = 1/(1+r)n

r=interest rate

n=number of times interest is paid.

2. PV Factor = 1/(1.1)2 = 0.8264

PV of the bond = 100*0.8264 = 82.64